# Mind the Gap: a Spectral Analysis of Rank Collapse and Signal Propagation in Attention Layers

### v2

@@ Abstract
Attention layers are the core component of transformers, the current state-of-the-art neural network architecture. Alternatives to softmax-based attention are being explored due to its tendency to hinder effective information flow. Even at initialisation, it remains poorly understood why the propagation of signals and gradients through these random networks can be pathological, resulting in issues known as (i) vanishing/exploding gradients and (ii) rank collapse in depth, i.e. when all tokens converge to a single representation along layers. While rank collapse in depth naturally arises from repeated matrix multiplications—a common pattern across various architectures—we identify an additional and previously unknown challenge unique to softmax attention layers: (iii) rank collapse in width, which occurs as the context length increases. Using Random Matrix Theory, we conduct a rigorous analysis that uncovers a spectral gap between the two largest singular values of the attention matrix as the cause of (iii), which in turn exacerbates (i) and (ii). Building on this insight, we propose a novel yet simple practical solution to mitigate rank collapse in width by removing the outlier eigenvalue(s). Our theoretical framework offers a fresh perspective on recent practical studies, such as [YDX+24, AGW23], whose ad hoc solutions can now be interpreted as implicit efforts to address the spectral gap issue. This work provides valuable theoretical support for ongoing large-scale empirical research, bringing theory and practice one step closer in the understanding of transformers. [^2] Our code is available at https://shorturl.at/0zk8q.
@@ 1.1 Related work
The rank collapse in transformers was first explored in [DCL21], where the authors show that the output of an attention-only transformer converges exponentially with depth to a single representation across tokens. [NAB+22] make a connection between rank collapse and vanishing gradients by assuming “uniform attention” $\mathbf{A}=\frac{1}{T}\mathbf{1}_{T\times T}$, essentially proving rank collapse in depth based on the assumption that rank collapse in width has already occurred (as opposed to showing why this premise holds).
@@ Appendix A Appendix
A.1 Proofs
@@ A.2 Lemmas
A.3 Can transformers achieve dynamical isometry?

### v3

@@ Abstract
Attention layers are the core component of transformers, the current state-of-the-art neural network architecture. Alternatives to softmax-based attention are being explored due to its tendency to hinder effective information flow. Even at initialisation, it remains poorly understood why the propagation of signals and gradients through these random networks can be pathological, resulting in issues known as (i) vanishing/exploding gradients and (ii) rank collapse in depth, i.e. when all tokens converge to a single representation along layers. While rank collapse in depth naturally arises from repeated matrix multiplications—a common pattern across various architectures—we identify an additional and previously unknown challenge unique to softmax attention layers: (iii) rank collapse in width, which occurs as the context length increases. Using Random Matrix Theory, we conduct a rigorous analysis that uncovers a spectral gap between the two largest singular values of the attention matrix as the cause of (iii), which in turn exacerbates (i) and (ii). Building on this insight, we propose a novel yet simple practical solution to mitigate rank collapse in width by removing the outlier eigenvalue(s). Our theoretical framework offers a fresh perspective on recent practical studies, such as [YDX+24, AGW23], whose ad hoc solutions can now be interpreted as implicit efforts to address the spectral gap issue. This work provides valuable theoretical support for ongoing large-scale empirical research, bringing theory and practice one step closer in the understanding of transformers. [^2] Our code is available at https://github.com/thizirinait/Spectral-Analysis-of-Attention-Layers.git.
@@ 1.1 Related work
Rank collapse in transformers was first explored in [DCL21], where the authors show that the output of an attention-only transformer converges exponentially with depth to a single representation across tokens. [NAB+22] make a connection between rank collapse and vanishing gradients by assuming “uniform attention” $\mathbf{A}=\frac{1}{T}\mathbf{1}_{T\times T}$, essentially proving rank collapse in depth based on the assumption that rank collapse in width has already occurred (as opposed to showing why this premise holds).
@@ 5 Conclusion
6 Acknowledgements
TNS is supported by the UK Engineering and Physical Sciences Research Council (EPSRC) through the grant EP/W523781/1. JT acknowledges support from His Majesty’s Government in the development of this research as well as by EPSRC grant EP/Y028872/1, Mathematical Foundations of Intelligence: An “Erlangen Programme” for AI.
@@ Appendix A Appendix
  • Section A.1 provides strong empirical support for the main theorem of this paper even when its assumptions are violated in practice.
@@
A.1 Justification of our theoretical assumptions
In this section, we empirically support the finding of our main theorem—a spectral gap arises in width within attention layers—even when its assumptions are violated. Our formal proof should be seen more as a proof of concept, and its conditions to hold are not necessary in practice, but rather made for mathematical rigor.
Inputs do not need to be isometric in practice.
In Figure 8, sentences are tokenised with a pre-trained tokeniser, making inputs to our randomly initialised BERT clearly non-isometric. We include $\mathbf{X}_{0}\mathbf{X}_{0}^{T}$ to illustrate its deviation from identity. When tokens are instead assigned random embeddings as in Figure 9, the covariance matrix is closer to orthogonality. Yet, the spectral gap and rank collapse persist in both cases. We also show the spectra of a random attention layer with (synthetic) non-isometric input tokens to mirror Figure 1 of the main in Figure 10.
(a) $\mathbf{X}_{0}\mathbf{X}_{0}^{\top}$ when $T=75$
(b) $\mathbf{X}_{0}\mathbf{X}_{0}^{\top}$ when $T=282$
(c) $\mathbf{X}_{0}\mathbf{X}_{0}^{\top}$ when $T=446$
(d) Eigenvalues in the complex plane of the attention matrix in the first head of layer $1$.
Figure 8: Spectral gap (second row) persists even when the inputs are not orthogonal (see first row). Experiment conducted with a randomly initialised RoBERTa encoder, hence $d=768$ and $T<512$ by construction, on real text data from our abstract, processed by a pretrained tokenizer available on HuggingFace. We present a single realisation, though this behavior is consistently observed across multiple runs.
(a) $\mathbf{X}_{0}\mathbf{X}_{0}^{\top}$ when $T=75$
(b) $\mathbf{X}_{0}\mathbf{X}_{0}^{\top}$ when $T=282$
(c) $\mathbf{X}_{0}\mathbf{X}_{0}^{\top}$ when $T=446$
(d) Eigenvalues of attention matrix in the first head of layer $1$ match our theoretical expectations.
Figure 9: Spectral gap (second row) persists on non-synthetic yet random data. Using a randomly initialized RoBERTa encoder ($d=768$, $T<512$ by construction), we substitute the word embeddings given by the pretrained tokenizer with random ones, making the input covariance matrix closer to identity (first row). We present a single realisation, though this behavior is consistently observed across multiple runs.
(a) $\gamma=0.1$
(b) $\gamma=0.5$
(c) $\gamma=2$
(d) $\gamma=10$
Figure 10: The eigenvalue distribution of a single attention layer for non-isometric synthetic input and different values of $\gamma=\frac{T}{d}$. The input is generated by drawing i.i.d. Gaussian entries followed by normalising the rows, so each token has unit length but they are not necessarily orthogonal. The gap between the bulk and the edge persists despite the bulk looking different for $\gamma>1$.
The context length $T$ does not need to be bigger than the token dimension $d$ in practice.
Note that without $T\leq d$, the inputs cannot be isometric so this extra assumption was solely made for this reason. We provide the reader with an ablation on the role of $\gamma$, the ratio between the context length and the token dimension. In Figure 10, we explore information propagation in TinyBert, for which $\gamma\geq 1$, hence necessarily the tokens must be non-orthonormal. The spectral gap persists, and rank collapse in width follows. Another ablation, this time conducted on synthetic data, can be found in Figure 10.
(a) Spectral gap in the attention matrix of the first head of layer $1$.
(b) Stable rank quickly collapses in width, even when $\gamma>1$.
Figure 11: Spectral gap (top) persists in TinyBert encoders for which $\gamma>1$, highlighting the generality of our result, which constrained the value of $\gamma$ only for the sake of having a mathematical proof. As an almost direct consequence of the spectral gap in the attention matrix, stable rank rapidly collapses (bottom) in width. Here $d=312$ by design and we increase $\gamma$ by feeding the network with longer sequences. Whilst the eigenvalues (top) are shown for one specific realisation (but consistently observed across runs), the stable rank (bottom) is averaged over $5$ runs.
A.2 Proofs
@@ A.3 Lemmas
A.4 Can transformers achieve dynamical isometry?